OpenStudy (anonymous):

Rewrite with only sin x and cos x. sin 3x - cos x

OpenStudy (cwrw238):

Hint sin2x = 2sinxcosx

OpenStudy (anonymous):

Can you help me? This confuses me too much

OpenStudy (cwrw238):

how far have you got with trigonometry?

OpenStudy (cwrw238):

sorry i have to go now

OpenStudy (anonymous):

do you know any trig identities for multiple of an angle?

OpenStudy (anonymous):

here is the one that is worth memorizing: sin(3x) = -4sin^3(x) + 3sin(x)

OpenStudy (anonymous):

Okay an you help me solve this

OpenStudy (anonymous):

@sourwing

OpenStudy (anonymous):

you're practically done. sin(3x) - cos(x) = sin(3x) = -4sin^3(x) + 3sin(x) - cos(x)

OpenStudy (anonymous):

A. 2 sin x - sin^3 x - cos x B. 2 sin x cos^2 x + sin x - 2 sin ^3 x - cos x C. 2 sin^3 x cos^4 x + 1 D. 3 sin x cos x - sin^3 x - cos x

OpenStudy (anonymous):

I think it's D @sourwing

OpenStudy (cwrw238):

sin3x = sin(2x + x) = sin2xcosx + sinxcos2x = 2sinxcos^2x + sinx( cos^2x - sin^2x) = 2sinxcos^2x + sinxcos^2x - sin^3x = 3sinx cos^2x - sin^3x sin3x - cosx = 3sinx cos^2x - sin^3x - cosx

OpenStudy (cwrw238):

- that result works though its not exactly = D